### v1

# Viewpoint Matters: Dynamically Optimizing Viewpoints with Masked Autoencoder for Visual Manipulation

Abstract
Robotic manipulation continues to be a challenge, and imitation learning (IL) enables robots to learn tasks from expert demonstrations. Current IL methods typically rely on fixed camera setups, where cameras are manually positioned in static locations, imposing significant limitations on adaptability and coverage. Inspired by human active perception, where humans dynamically adjust their viewpoint to capture the most relevant and least noisy information, we propose MAE-Select, a novel framework for active viewpoint selection in single-camera robotic systems. MAE-Select fully leverages pre-trained multi-view masked autoencoder representations and dynamically selects the next most informative viewpoint at each time chunk without requiring labeled viewpoints. Extensive experiments demonstrate that MAE-Select improves the capabilities of single-camera systems and, in some cases, even surpasses multi-camera setups. The project will be available at https://mae-select.github.io.
Index Terms—  Manipulation, Imitation Learning, Active Perception
1 Introduction
Robotic manipulation is a core challenge in robotics. Imitation Learning (IL) [17, 12, 6, 5, 2] has become a leading approach for enabling robots to learn complex tasks through expert demonstrations. Recently, advances in deep generative models [7, 10, 14, 9], have empowered IL by allowing robots to process high-dimensional sensory inputs, leading to promising results in robotic manipulation [3, 25, 1, 19].
However, most current IL methods rely on fixed camera setups, either single or multiple cameras, which pose significant limitations. In fixed single-camera setups [16, 18], though cost-effective, robots face challenges due to the limited field of view, which may obstruct critical parts of the environments, negatively impacting task performance. Multi-camera setups, while designed to provide more comprehensive scene coverage, introduce their own complexities: the abundance of redundant or irrelevant information can overwhelm learning algorithms and decrease efficiency. As shown in Sec.3, these passive static multi-view setups do not always provide the most task-relevant and cleanest information, leading to suboptimal decision-making.
In contrast, humans dynamically adjust their viewpoints while performing tasks. By actively moving our head, we naturally seek the most informative, least noisy perspectives. Inspired by this human capability, we propose shifting from passive, static perception to active perception, where the viewpoint is dynamically adjusted throughout the task to optimize information intake. In a practical robotic setting, this could be embodied by a robot moving its head to capture the most task-relevant views in real-time. In this paper, we focus on the feasibility of active viewpoint selection for robotic manipulation as an initial exploration along this direction.
To this end, we introduce MAE-Select, a framework designed to actively select optimal viewpoints for single-camera robotic setups. MAE-Select first fully utilizes the powerful pre-trained representations from the multi-view masked autoencoders (MAEs) [8, 23], leveraging its complete encoder-decoder architecture to obtain estimated multi-view representations. Unlike prior works that focus on fixed viewpoints [21], MAE-Select dynamically predicts the next better viewpoint based on the current chunk of visual and action information. Crucially, this viewpoint selection is learned solely through imitation learning, which does not require manual labels for optimal views. It demonstrates significant potential in the advancement of single-camera robotic manipulation.
Fig. 1: Illustration of our proposed method. Left depicts the pre-training stage of the multi-view masked autoencoder with joint embeddings. Middle illustrates the training process of our framework using imitation learning. Right demonstrates how the framework operates during inference.
Our key contributions are as follows:
  • We propose MAE-Select, a novel mechanism that dynamically selects the next optimal viewpoint at each time chunk without manual labels.
  • We present an IL framework that fully exploits pre-trained MAE representations for manipulation.
  • We demonstrate through various experiments that MAE-Select enhances manipulation accuracy in single-camera setups, even outperforming multi-camera systems in certain tasks.
2 Method
We introduce MAE-Select, a novel framework that enables robotic agents to actively select informative viewpoints for manipulation tasks. The core of our approach is a new formulation in which optimal viewpoint selection is learned implicitly through an imitation learning objective, without requiring explicit supervision or reinforcement learning. Unlike previous work that only uses pre-trained encoders [20, 21], our method exploits both encoder and decoder of a multiview masked autoencoder (MV-MAE), enabling the agent to construct a rich scene representation from a single view, crucial for accurate action prediction and informed viewpoint selection. An overview of our architecture is shown in Fig.1.
2.1 Problem Formulation
We frame the robotic manipulation task as a learning problem over a sequence of observations and actions. Let the agent’s state at time $t$ be composed of its proprioceptive information $s_{t}$ (e.g., joint angles) and a set of visual observations $O_{t}=\{o_{t}^{v}\}_{v\in V}$ from a fixed set of $N_{v}$ available viewpoints $V$. The agent’s goal is to learn a policy $\pi$ that takes the current state and generates a chunk of $T$ future actions, $a_{t:t+T-1}=\{a_{t},\dots,a_{t+T-1}\}$.
In our active perception setting, the policy does not have access to all viewpoints simultaneously. Instead, it operates on a single selected viewpoint $o_{t}^{v_{t}}$ at time $t$. Therefore, the objective is to learn two key policies:
  1. 1.
    An action policy $\pi_{\theta}(a_{t:t+T-1}|o_{t}^{v_{t}},s_{t})$: This policy predicts a sequence of future actions based on the current single-view observation and proprioceptive state.
  2. 2.
    A view selection policy $\pi_{\psi}(v_{t+T}|o_{t}^{v_{t}},s_{t},a_{t:t+T-1})$: This policy selects the most advantageous viewpoint, $v_{t+T}$, for the subsequent time chunk, conditioned on the information from the current chunk.
Both policies, parameterized by $\theta$ and $\psi$ respectively, are trained jointly from a dataset of expert demonstrations $\mathcal{D}=\{(O_{t},s_{t},a_{t})\}_{t=0\dots T_{m}}^{N}$ using an imitation learning framework.
2.2 Multi-View Masked Autoencoder Pre-training
We pre-train a Multi-View Masked Autoencoder (MV-MAE) on demonstration data to learn a compact, 3D-aware scene representation (Fig. 1, left).
Given a multi-view observation $O_{t}$, we extract feature maps $\{\bar{o}_{t}^{v}\}_{v\in V}$ and apply dual masking: (1) patch masking, which randomly masks a large portion of patches within each view, and (2) view masking, which randomly drops entire views to encourage cross-view reasoning. The remaining patches, with view and position embeddings, are fed into a Transformer encoder $f_{\phi}$ to produce latent $z_{t}^{m}$. A decoder $g_{\phi}$ reconstructs all views from $z_{t}^{m}$, mask tokens, and the robot state embedding $s_{t}$ by minimizing
$\mathcal{L}_{\text{MAE}}=\mathbb{E}_{O_{t},s_{t}}\left[\|O_{t}-g_{\phi}(f_{\phi}(O_{t}^{m}),s_{t})\|^{2}_{2}\right].$ (1)
This pre-training enables the model to infer full 3D scenes from partial or occluded inputs for downstream tasks.
2.3 Next Better Viewpoint Selection
Following pre-training, we jointly fine-tune the model and train both the action and viewpoint selection policies (Fig. 1, middle). Although action decoder training uses ground-truth actions, defining a ground-truth “optimal” viewpoint is impractical, as it depends on task and context. Instead of reinforcement learning, which is difficult due to the sim-to-real gap, we propose a label-free strategy to train the viewpoint selector $\pi_{\psi}$. Specifically, the action prediction loss on a future data chunk $\mathcal{D}_{t+T}$ is used as a supervisory signal for the viewpoint chosen in the current chunk $\mathcal{D}_{t}$. We apply this process to two consecutive trajectory chunks, $\mathcal{D}_{t}$ and $\mathcal{D}_{t+T}$.
Processing the Current Chunk ($\mathcal{D}_{t}$): For the current chunk starting at time $t$, we begin with a single view $o_{t}^{v}$, selected uniformly at random. This single-view observation is passed through the entire pre-trained MV-MAE ($f_{\phi},g_{\phi}$) to generate an estimated multi-view feature context $C_{t}=g_{\phi}(f_{\phi}(o_{t}^{v}),s_{t})$. This context $C_{t}$ is then fed into the action decoder $\pi_{\theta}$, a diffusion-based policy, which is trained to predict the noise $\epsilon_{t}$ added to the expert action trajectory $a_{t:t+T-1}$. The action loss for this chunk is computed as:
$\mathcal{L}_{\text{action}}^{(t)}=\mathbb{E}_{\epsilon_{t},k}\left[\|\epsilon_{t}-\pi_{\theta}(C_{t},a_{t:t+T-1}^{0}+\epsilon_{t},k)\|^{2}\right]$ (2)
where $k\sim\mathcal{U}(1,K)$ is a randomly sampled diffusion timestep.
Selecting the View for the Next Chunk ($\mathcal{D}_{t+T}$): The viewpoint selector $\pi_{\psi}$, implemented as a Transformer encoder, takes the feature context $C_{t}$ and the ground-truth action trajectory $a_{t:t+T-1}$ from the current chunk to predict a probability distribution of viewpoints for the next chunk:
$\hat{c}=\text{Softmax}(\pi_{\psi}(C_{t},a_{t:t+T-1}))$ (3)
To select a viewpoint while maintaining differentiability for backpropagation, we employ a straight-through estimator as VQ-VAE [24]. In the forward, we perform an argmax operation to obtain one-hot vector $\mathbf{y}$ representing the chosen view. For the backward, gradients flow through the continuous softmax probabilities $\hat{c}$. This is implemented as:
[…]
where $\text{sg}[\cdot]$ denotes the stop-gradient operator. This mechanism enables the gradient from the subsequent action loss to inform the selection policy.
Processing the Next Chunk ($\mathcal{D}_{t+T}$): The one-hot vector $\mathbf{y}$ is used to select a single viewpoint $o_{t+T}^{\hat{v}}$ from the next observation set $O_{t+T}$. This selection is performed via a dot product $o_{t+T}^{\hat{v}}=\mathbf{y}\cdot O_{t+T}$, which also enables gradient flow back to the selection policy. This selected view is then processed through the identical pipeline as the first chunk, yielding a corresponding action loss, $\mathcal{L}_{\text{action}}^{(t+T)}$.
Updating the Viewpoint Selector: There is no explicit “view loss.” The viewpoint selector $\pi_{\psi}$ is trained using gradients from the future action loss $\mathcal{L}_{\text{action}}^{(t+T)}$, backpropagated through the action decoder and selected observation via STE, directly optimizing $\pi_{\psi}$ to select viewpoints that minimize future action prediction error.
The overall objective combines both action losses and the auxiliary reconstruction loss:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{action}}^{(t)}+\lambda_{1}\mathcal{L}_{\text{action}}^{(t+T)}+\lambda_{2}\mathcal{L}_{\text{MAE}}$ (5)
where $\pi_{\theta}$ is updated by both action losses, while $\pi_{\psi}$ is optimized only by the future action loss.
Inference: During inference (Fig. 1, right), the process is autoregressive. Starting from a random initial viewpoint, it predicts both the current action chunk and the optimal viewpoint for the next chunk. The predicted action is then used alongside the newly selected viewpoint for the subsequent step, creating a dynamic perception-action loop.
Table 1: Results of comparison experiment. * represents with Disturbance. For the eight simulated tasks, we report the success rate, with 50 policy evaluations each. For the real-world tasks, we report the number of successes, training with human data, with 10 evaluations. For Diffusion Policy and MAE-Diffusion, the viewpoint data used for training is the same as the viewpoint used for testing. For MAE-Select, we use both viewpoints for training and only one viewpoint is used per time chunk during testing. Bold and underlined fonts mean the best and second-best results.
Method Bimanual Insertion Put Box In Cabinet Put Box In Bin Put Box In Bin*
Top Front Both Top Left Both Top Left Both Top Left Both
Diffusion Policy [3] 42% 44% 50% 16% 18% 26% 80% 64% 84% 38% 30% 44%
MAE-Diffusion 48% 50% 54% 42% 42% 46% 84% 78% 92% 52% 46% 60%
MAE-Select 52% 50% 88% 58%
Method Phone On Base Pick Up Cup Take Umbrella Unplug Charger
Front Wrist Both Front Wrist Both Front Wrist Both Top Wrist Both
Diffusion Policy [3] 82% 56% 78% 60% 40% 64% 58% 36% 54% 44% 30% 34%
MAE-Diffusion 86% 70% 88% 68% 66% 62% 56% 42% 64% 46% 34% 52%
MAE-Select 92% 70% 60% 58%
Method Put Eggplant To Bowl Put Eggplant To Bowl* Put Bitter Melon In Cabinet
Top Wrist Both Top Wrist Both Top Left Both
Diffusion Policy [3] 2/10 1/10 5/10 2/10 0/10 3/10 0/10 1/10 2/10
MAE-Diffusion 4/10 4/10 7/10 4/10 3/10 6/10 2/10 4/10 4/10
MAE-Select 6/10 6/10 5/10
Fig. 2: Visualization of the selected viewpoints in our experiments, showcasing both simulation and real-world environments. Each row represents the procedure of a specific task, indicating the necessity of selecting different viewpoints.
3 Experiments
We evaluate MAE-Select across 3 challenging scenarios and 8 demanding tasks in simulation, including simulations in ACT [25], RLBench [11], and our designed robot in MuJoCo [22], along with three real-world tasks.
3.1 Implementation
Table 2: Results of ACT experiments. We report the success rate, with 50 policy evaluations each. Bold and underlined fonts mean the best and second-best results.
Method Bimanual Insertion Phone On Base
Top Front Both Front Wrist Both
ACT 14% 26% 34% 56% 50% 58%
MAE-ACT 28% 30% 42% 60% 58% 66%
MAE-Select 36% 70%
Table 3: Ablation Studies on MAE Encoder and Decoder Utilization.
Method Put Box In Cabinet Phone On Base
Top Left Both Front Wrist Both
MAE-Encoder 20% 28% 34% 76% 56% 80%
MAE-Diffusion 42% 42% 46% 86% 70% 88%
We explore a single-camera control setup where the system is trained with multiple camera views but deployed with a single moving camera. At each time chunk, the camera is placed at one of the training viewpoints, enabling flexible operation in real-world settings where only one camera is available. We base our implementation on the architecture of the diffusion policy [3]. The action space corresponds to the joint angles of the robot arm, while the image observations have a resolution of $224\times 224$ with a patch size of 16. Our masked autoencoder utilizes a 12-layer ViT [4] for encoder and an 8-layer ViT for decoder, with an embedding dimension of 512. During pretraining, we use a batch size of 128 over 100 epochs. For RLBench and real-world tasks, the time chunk size is set to 20, with a total of 600 epochs. For other tasks, the time chunk size is 100, with 1,000 epochs in total. $\lambda_{1}$ and $\lambda_{2}$ are empirically set to 2.0 and 1.0, respectively. For each method, we evaluate the best-performing checkpoints from the last three evaluated at 100-intervals, with 50 environment initializations in simulation. All models were trained and tested on NVIDIA RTX 4090 GPUs.
3.2 Results
We compare MAE-Select with Diffusion Policy [3] and its MAE-enhanced variant (MAE-Diffusion) across various tasks. Unlike the baselines, which use fixed viewpoints during training and testing, MAE-Select leverages multi-view training and dynamically selects the most informative single viewpoint for the next time chunk during inference.
As shown in Tab.1, MAE-Select consistently outperforms fixed single-camera setups in both simulation and real-world experiments—for instance, achieving an 8% gain over the best fixed camera method and 32% over prior work in the Put Box In Cabinet task. Its strength lies in intelligently exploiting limited visual input through optimal viewpoint selection. Interestingly, in some tasks a single camera even surpasses multi-camera setups (e.g., Diffusion Policy on Unplug Charger: top view 44% vs. two views 34%), likely due to noise and alignment issues from multiple inputs. MAE-Select avoids such pitfalls, remaining competitive and often superior, even compared with multi-camera configurations.
To analyze system behavior under single-camera control, we visualize deployment phases (Fig.2). MAE-Select adaptively selects optimal viewpoints, focusing on critical regions while ignoring irrelevant areas.
3.3 Ablation Study
Our viewpoint selection method is agnostic to the choice of action decoder, as evidenced by its integration with ACT [25] (Tab.2), confirming its adaptability across architectures. We further conduct an ablation study comparing the full encoder-decoder masked autoencoder with a variant using only the encoder [20, 21]. Results (Tab.3) show that leveraging the full structure significantly boosts performance, especially under occlusion, highlighting the decoder’s role in refining visual representations and improving generalization in both simulation and real-world tasks.
4 Conclusions and Future Work
We propose MAE-Select, a framework that enables single-camera robots to actively optimize viewpoints for manipulation tasks. By exploiting pre-trained multi-view masked autoencoder representations and dynamically selecting informative viewpoints at each time chunk without manual annotations, MAE-Select boosts manipulation efficiency and can even outperform multi-camera systems. Its main limitation is optimizing over discrete rather than continuous viewpoints, which limits flexibility in dynamic environments. Future work could integrate NeRF [15] or 3D Gaussian processes [13] for continuous viewpoint optimization.